OpenStudy (anonymous):

find the equation of the axis of symmetry and the vertex of y= -3x^2 + 8x +7 and y= 2x^2 +4x-3

OpenStudy (anonymous):

x=-b/2a is the equation for the axis of symmetry. so -8/6=-4/3. x=-4/3 is the equation of the axis of symmetry. Plugging that x-value back into the first equation gives (-4/3, and the value you get when you plug -4/3 back into the equation). For the second one, -b/2a=-4/4=-1. x=-1 is the equation of the axis of symmetry. Plug that value back in to get (-1, y)

OpenStudy (anonymous):

thank you!! :)